# A Unified Framework for Direction and Diffuseness Estimation Using Tight-Frame Microphone Arrays

### v2

@@ Abstract
Submitted to the Journal of the Acoustical Society of America (under review).
@@ I Introduction
A diffuse sound field is an ideal state in which both sound pressure and particle velocity are spatially and directionally isotropic. This assumption underpins the classical reverberation theories of Sabine and Eyring [1, 2, 3] and remains central in architectural acoustics [4, 5]. Quantitative evaluation of “diffuseness” thus plays a key role—not only for objectively assessing sound fields but also in applications such as room design, spatial reproduction, and auditory analysis.
@@
In parallel, Thiele visualized reflection directions and strengths in concert halls, linking diffuseness with perceptual attributes such as clarity [11, 12, 13]. Following this, Gover and colleagues [14, 15] developed spherical microphone-array systems capable of analyzing the directional and spatial variations of reverberant fields. By steering beams over multiple directions, they computed the variance of directional power to define “directional diffusion” and the anisotropy index, showing how diffuseness evolves over time as the reverberant field becomes more isotropic and later anisotropic. These studies established the idea that a diffuse field can be characterized by the uniformity of directional energy distribution, forming an experimental foundation for later diffuseness metrics.
From the late 1980s to 1990s, the energy-transport equation provided a physically grounded framework for describing energy propagation in rooms. Derived from Boltzmann’s transport theory, it expresses temporal variation of acoustic energy density as a balance between flux and loss terms. Applying a Fourier-type approximation to the flux term yields the acoustic diffusion equation, which models energy distribution continuously in enclosed spaces. Picaut et al. [16] developed a numerical model based on this equation, later extended by Valeau et al. [17] to include anisotropy and boundary conditions. Hanyu [18] further incorporated wall scattering and absorption coefficients to characterize room diffuseness quantitatively. These developments form the theoretical core of modern room-diffusion analysis. Following this line of study, Sakuma [19] and Omoto [20] extended the framework toward more realistic nondiffuse conditions, introducing analytical models that link the statistical treatment of scattering with the observable decay characteristics in rectangular rooms. Cox and D’Antonio [21] organized diffuser design and diffuse-field modeling into practical methodologies, and the ISO 17497 series [22, 23] standardized measurement and evaluation methods internationally.
In the 2000s, diffuseness quantification in time and space advanced rapidly. The First Order Ambisonics (FOA) based intensity-energy ratio (I/E) method by Merimaa and Pulkki [24, 25] characterized and reproduced sound fields by their diffuseness. Acoustic intensity was obtained from impulse responses, allowing estimation of diffuseness and major reflection directions, later reproduced using Vector Base Amplitude Panning (VBAP) [26] in Spatial Impulse Response Rendering (SIRR). The method was extended to continuous signals as Directional Audio Coding (DirAC) [27, 28], maintaining the same intensity-diffuseness principle. Since then, the method has evolved to estimate prominent reflections from the generalized cross-correlation of signals acquired with multiple microphones, named Spatial Decomposition Method, (SDM) [29], and more recently, applications in the ultrasonic domain [30] have also been reported.
@@
Following the covariance-based formulation by Epain and Jin[32], Politis and Pulkki[34, 33] extended this concept to spatially weighted energy analysis and parametric spatial audio frameworks. Their work unified the estimation of intensity, energy density, and diffuseness in the HOA domain and demonstrated its applicability to spaced microphone arrays and multichannel rendering. More recently, McCormack et al. [35] investigated the perceptual effects of spherical order, dedicated diffuse rendering, and frequency resolution in higher-order Spatial Impulse Response Rendering (SIRR), highlighting that perceived spaciousness strongly depends on the accuracy of diffuse-field modeling.
Throughout this history, the link between diffuseness estimation and intensity measurement has been widely recognized, as clearly demonstrated by the I/E method. The p-p probe method summarized by Fahy [36, 37] estimates local energy flow and reflection distributions. Guy and Abdou [38] visualized reflections and energy flow, while de Bree et al. [39, 40] introduced pressulre-velocity sensors. Zhang et al. [41] applied this p-u probes to near-field acoustic holography.
The principle of acoustic intensity measurement using oppositely oriented cardioid microphones was first proposed by Bauer [42], forming the basis of what is now known as the C-C method. This approach was later refined and systematically established by Hoshi and Hanyu [43], also Hanyu and Hoshi [44], who developed reliable formulations for practical intensity measurements in various acoustic environments. These studies advanced an integrated view of intensity and diffuseness, linking energetic analysis with spatial measurement.
@@
While the proposed framework enables stable and broadband evaluation, it also reveals a fundamental limitation of conventional approaches: Conventional methods often reduce diffuseness to a single scalar, neglecting directional variation. In real environments, however, diffuseness depends on surface reflection and absorption distributions, making directional diffuseness essential. Its prototype appears in Meyer’s “hedgehog” visualization of directional energy using a parabolic microphone [12, 13]. This approach evolved into the SDM [29], psychoacoustic studies by Amengual et al. [45], and visualization work by Pätynen et al. [46], forming the basis for direction-dependent diffuseness analysis.
Building upon these theoretical foundations, the present study extends diffuseness estimation to practical microphone arrays with several different perspectives:
Microphone array based on tight-frame theory and spatial sampling — A tight frame is an overcomplete basis that enables redundant yet stable reconstruction [47]. On the sphere, a spherical $(t,t)$-design with $t=1$ is equivalent to a tight frame [48], providing a foundation for isotropic sampling in spherical processing. Recent work extends this concept to spherical framelets [49], optimal finite-dimensional frames[50]. Building on these, the present study implements uniform directional sampling based on a tight-frame array for diffuseness and intensity measurement.
@@
Systematic comparison across indices and array configurations — Three array types: A-format, Tight Frame, and Rigid Sphere — are compared under identical conditions to evaluate consistency, frequency dependence, and the effects of spherical-harmonic order and whitening. Emphasis is placed on the trade-off between the numerical instability of high-order spherical harmonic (HOA) analysis [32] and the implementation stability of first-order (FOA) approaches. Overall, these analyses aim to bridge theoretical validity and practical reproducibility, establishing a unified framework for diffuseness evaluation.
@@ II.1 A-format microphone (Afmt)
where $Z_{0}$ is the characteristic acoustic impedance of the medium. From these quantities, both intensity and the total energy can be derived as
@@ II.3 Tight frame microphone array (TF24)
We propose a new array using multiple highly directional microphones arranged isotropically on a sphere based on a tight-frame distribution, referred to as TF24. A finite set of such points uniformly approximates the sphere while maintaining reconstruction stability. This microphone configuration was initially devised for measuring directional impulse responses intended for sound field reproduction[51]. Preliminary trials were conducted to estimate acoustic intensity using this setup, which subsequently led to the more detailed investigation presented in this study.
@@
$\bm{R}=\begin{bmatrix}0&1/\sqrt{2}&1/\sqrt{2}\\ 1/2&1/2&1/\sqrt{2}\\ 1/\sqrt{2}&0&1/\sqrt{2}\\ 1/2&-1/2&1/\sqrt{2}\\ 0&-1/\sqrt{2}&1/\sqrt{2}\\ -1/2&-1/2&1/\sqrt{2}\\ -1/\sqrt{2}&0&1/\sqrt{2}\\ -1/2&1/2&1/\sqrt{2}\\ 0&1&0\\ 1/\sqrt{2}&1/\sqrt{2}&0\\ 1&0&0\\ 1/\sqrt{2}&-1/\sqrt{2}&0\end{bmatrix}$, (13)
@@
Microphone pairs aligned along twelve directions compute pseudo-intensity via the cardioid-pair (C-C) principle [43, 44], here adapted for narrow-directional microphones as the pseudo C-C method.
@@
When two microphones are aligned along the same axis but oriented in opposite directions, their directional responses can be expressed as $d(\theta)$ and $d(\pi-\theta)$, respectively. The sum and difference of their complex amplitudes,
$M^{+}_{\bm{r}_{i}}\pm M^{-}_{\bm{r}_{i}},$
then form symmetric and antisymmetric combinations of these directional responses. The sum component primarily exhibits an omnidirectional characteristic, approximating the local sound pressure, whereas the difference component resembles a figure-of-eight pattern, corresponding to the particle velocity component along $\bm{r}_{i}$. Hence, the pairwise sum and difference of oppositely oriented directional microphones provide a practical means of approximating the local pressure and the corresponding particle-velocity component.
@@
$p_{\bm{r}_{i}}$ $=M^{+}_{\bm{r}_{i}}+M^{-}_{\bm{r}_{i}},$ (15)
$u_{\bm{r}_{i}}$ $=\frac{M^{+}_{\bm{r}_{i}}-M^{-}_{\bm{r}_{i}}}{Z_{0}},$ (16)
$I_{\bm{r}_{i}}$ $=\frac{|M^{+}_{\bm{r}_{i}}|^{2}-|M^{-}_{\bm{r}_{i}}|^{2}}{2Z_{0}},$ (17)
$E_{\bm{r}_{i}}$ $=\frac{|p_{\bm{r}_{i}}|^{2}}{4\rho_{0}c^{2}}+\frac{\rho_{0}}{4}|u_{\bm{r}_{i}}|^{2}.$ (18)
Figure 3: Schematic diagram of the signals used to construct the pseudo C–C method in the TF24 system. Twelve directional pairs are formed from 24 microphones.
In practice, the directional elements used in the array do not have strictly cardioid directivity but rather exhibit sharper lobes depending on frequency. Therefore, the pressure obtained from their summation and the velocity inferred from their differential output cannot be regarded as the exact physical quantities. Instead, they represent operational approximations of the local sound pressure and particle velocity, derived under the assumption of complementary directivity pairs. The corresponding acoustic intensity and energy density are thus defined in this approximate sense and referred to here as pseudo-intensity and pseudo-energy density. Representative examples of these directivity patterns are presented in the next section.
Similarly, the directional pseudo-diffuseness index based on the I/E concept is defined as
$\Psi_{\bm{r}_{i}}=1-\frac{|I_{\bm{r}_{i}}|}{E_{\bm{r}_{i}}c},$ (19)
@@ III.1 Overview of simulation framework
The Afmt array was modeled after the AMBEO VR Mic, and the capsule directivities were approximated by a cardioid model consistent with DPA 2012. For the TF24 array, we used a highly directional model consistent with DPA 4017. In both cases, directivity patterns were measured in an anechoic chamber at $3^{\circ}$ or $5^{\circ}$ resolution and fitted as a polynomial in $\cos\theta$ (power-series form) up to 8th order; these fits were used in simulation (Fig. 4).
Figure 4: Measured (solid lines) and modeled (broken lines) directivities of the (a) DPA2012 cardioid microphone assumed in Afmt array and (b) DPA 4017 narrow-directional microphone used in the TF24 array. To avoid visual clutter, the polar plots show the responses for 250 Hz to 1 kHz on the left half and for 2 kHz to 8 kHz on the right half. In the calculations presented in the text, however, a wider frequency range from 63 Hz to 16 kHz was analyzed in octave bands.
Figure 5: Example of modeled directional responses derived from the microphone directivity polynomial of Eq. (21) at 1 kHz. The outer line shows the pseudo-omnidirectional response (sum signal), and the inner line shows the pseudo-figure-of-eight response (difference signal). Both were obtained from paired microphones placed in opposite directions in the TF24 array.
Array geometries were set as follows (see Fig. 2): the A-format capsules are located 6 mm from the origin [Fig. 2(a)]; the rigid-sphere array uses a radius of 42 mm [Fig. 2(b)]. The spacing between microphones in the TF24 array was set to 20 mm; that is, the distance from the origin in [Fig. 2(c)] is 10 mm. Although this dimension may seem unrealistic for highly directional microphones approximately 200 mm in length, it represents a practical configuration assuming sequential impulse-response measurements by mechanically rotating a single microphone.
@@
For the present simulations, each directivity pattern was modeled as a polynomial function of $\cos\theta$, where $\theta$ is the angle between $\hat{d}_{m}$ and $\hat{\nu}$:
@@
As an example, Fig. 5 illustrates the modeled directivities obtained from the sum and difference of opposite microphone pairs based on the polynomial model in Eq. (21) at 1 kHz. The sum response approximates a pseudo-omnidirectional pattern, while the difference response approximates a pseudo-figure-of-eight pattern. For the Fibo64 rigid-sphere model with omnidirectional elements, $D_{m}=1$; directivity/baffle effects are captured by the radial filters $b_{n}(ka)$ in the Spherical Harmonic expansion. Thus, after computing $p_{m}(f)$ via (20), HOA coefficients are estimated from the SH model as described in Sec. II.
@@ III.2 Benchmark case definitions
A single plane wave arrives from a given direction; the theoretical diffuseness is $\Psi=0$. We sweep both elevation and azimuth over the sphere to test directional robustness and verify how accurately intensity-based direction-of-arrival can be recovered. For Case 1, 72 azimuthal directions (0 to 355) and 35 zenith angles (5 to 175), both at 5 intervals, were simulated, resulting in 2520 incidence directions in total. For each case, the mean diffuseness and the error in arrival-angle detection based on the acoustic intensity were evaluated to assess the accuracy of sound-arrival direction estimation for each array.
@@
In this case, an interference field was simulated by superposing two plane waves: a primary wave incident from the zenith direction and a secondary wave arriving from variable zenith angles between 0° and 180°. For each configuration, 1,000 realizations were generated by randomly varying the phase (within $\pm 2\pi$) and amplitude (within $\pm 3$ dB) of the secondary wave. The acoustic intensity, energy density, and covariance matrix were computed for each realization, and the final values were obtained by averaging these quantities over all realizations to represent the mean behavior of the interference field.

### v3

# A framework for diffuseness evaluation using a tight-frame microphone array configuration

@@ I Introduction
A diffuse sound field is an ideal state in which both sound pressure and particle velocity are spatially and directionally isotropic. This assumption underpins the classical reverberation theories and remains central in architectural acoustics. Quantitative evaluation of “diffuseness” thus plays a key role—not only for objectively assessing sound fields but also in applications such as room design, spatial reproduction, and auditory analysis.
@@
In parallel, Thiele visualized reflection directions and strengths in concert halls, linking diffuseness with perceptual attributes such as clarity [6, 7, 8]. Following this, Gover and colleagues [9, 10] developed spherical microphone-array systems capable of analyzing the directional and spatial variations of reverberant fields. By steering beams over multiple directions, they introduced anisotropy and directional diffusion measures based on the variance of directional power, establishing the idea that diffuseness can be characterized by the uniformity of directional energy distribution.
In the 2000s, diffuseness quantification in time and space advanced rapidly. The First Order Ambisonics (FOA) based intensity-energy ratio (I/E) method by Merimaa and Pulkki [11, 12] characterized sound fields through the relationship between acoustic intensity and energy density. Acoustic intensity was obtained from impulse responses, allowing the estimation of diffuseness and major reflection directions, later reproduced using Vector Base Amplitude Panning (VBAP)[13] in Spatial Impulse Response Rendering (SIRR). The same principle was extended to continuous signals in Directional Audio Coding (DirAC)[14, 15].
In parallel, diffuseness measures based directly on the statistical properties of acoustic intensity vectors were proposed. In particular, Ahonen and later Del Galdo et al. introduced indicators derived from the coefficient of variation of the intensity magnitude, relating diffuseness to the ratio between the norm of the mean intensity vector and the expected instantaneous intensity [16, 17].
@@
Following the covariance-based formulation by Epain and Jin[19], Politis and Pulkki[21, 20] extended this concept to spatially weighted energy analysis and parametric spatial audio frameworks. Their work unified the estimation of intensity, energy density, and diffuseness in the HOA domain. More recently, McCormack et al. [22] investigated the perceptual effects of spherical order, dedicated diffuse rendering, and frequency resolution in higher-order Spatial Impulse Response Rendering (SIRR), highlighting that perceived spaciousness strongly depends on the accuracy of diffuse-field modeling.
Throughout this history, the link between diffuseness estimation and intensity measurement has been widely recognized, as clearly demonstrated by the I/E method.
In addition to the conventional pressure-pressure method, the principle of acoustic intensity measurement using oppositely oriented cardioid microphones was first proposed by Bauer [23], forming the basis of what is now known as the C-C method. This approach was later refined and systematically established by Hoshi and Hanyu [24], also Hanyu and Hoshi [25], who developed reliable formulations for practical intensity measurements in various acoustic environments. These studies advanced an integrated view of intensity and diffuseness, linking energetic analysis with spatial measurement.
@@
Building upon these theoretical foundations, the present study extends diffuseness estimation to practical microphone arrays with the following specific contributions:
Microphone array based on tight-frame theory and spatial sampling — A tight frame provides a redundant yet stable sampling strategy on the sphere, allowing uniform directional coverage without requiring exact spherical harmonic reconstruction. We propose a new microphone array shape based on this theory.
@@
Systematic comparison across indices and array configurations — Three array types: A-format, Tight Frame, and Rigid Sphere — are compared under identical conditions to evaluate consistency, frequency dependence, and the effects of spherical-harmonic order and whitening. Emphasis is placed on the trade-off between the numerical instability of high-order spherical harmonic (HOA) analysis [19] and the implementation stability of first-order (FOA) approaches. Overall, these analyses aim to bridge theoretical validity and practical reproducibility, thereby establishing a consistent, practically applicable framework for diffuseness evaluation.
@@ II.1 A-format microphone (Afmt)
where $Z_{0}$ is the characteristic impedance of the medium. From these quantities, both intensity and the total energy can be derived as
@@ II.3 Tight frame microphone array (TF24)
A tight frame is an overcomplete basis that enables redundant yet stable reconstruction [26]. On the sphere, a spherical $t$-design with $t=1$ is equivalent to a tight frame [27], providing a foundation for isotropic sampling in spherical processing. Recent work extends this concept to spherical framelets [28], optimal finite-dimensional frames[29]. Building on these, the present study implements uniform directional sampling based on a tight-frame array for diffuseness and intensity measurement.
In practice, we propose a new array using multiple directional microphones arranged isotropically on a sphere, based on a tight-frame distribution, referred to as TF24. A finite set of such points uniformly approximates the sphere while maintaining reconstruction stability. This microphone configuration was initially devised for measuring directional impulse responses intended for sound field reproduction[30]. Preliminary trials were conducted to estimate acoustic intensity using this setup, which subsequently led to the more detailed investigation presented in this study.
@@
$\bm{R}=\begin{bmatrix}1/\sqrt{2}&0&1/\sqrt{2}\\ 1/2&1/2&1/\sqrt{2}\\ 0&1/\sqrt{2}&1/\sqrt{2}\\ -1/2&1/2&1/\sqrt{2}\\ -1/\sqrt{2}&0&1/\sqrt{2}\\ -1/2&-1/2&1/\sqrt{2}\\ 0&-1/\sqrt{2}&1/\sqrt{2}\\ 1/2&-1/2&1/\sqrt{2}\\ 1&0&0\\ 1/\sqrt{2}&1/\sqrt{2}&0\\ 0&1&0\\ -1/\sqrt{2}&1/\sqrt{2}&0\end{bmatrix}$, (13)
@@
As descrived above, the directional configuration of the Tight Frame array, based on vectors separated by 45 in azimuth and elevation, was originally devised by the authors as a practical microphone geometry for a hedgehog-type array intended for sound-field reproduction[30]. Microphones oriented exactly toward the zenith and the nadir were intentionally omitted, primarily for reasons of mechanical simplicity and ease of installation and handling. As a result, the adopted configuration exhibited more stable numerical properties.
For example, if the zenith and nadir directions are additionally included, resulting in a 14×3 directional matrix in Eq. (13), the matrix product $\bm{R}^{\!\top}\bm{R}$ no longer reduces to a scalar multiple of the identity matrix. In such a case, the physical quantities cannot be consistently transformed into the Cartesian $x$, $y$, and $z$ components using Eq. (9), thereby undermining the tight-frame property required for stable processing.
@@
Microphone pairs aligned along twelve directions compute pseudo-intensity via the cardioid-pair (C-C) principle [24, 25]. When measured microphone directivities are used, the corresponding quantities are distinguished by the prefix “pseudo-”, hereafter.
@@
When two microphones are aligned along the same axis but oriented in opposite directions, their directional responses can be expressed as $d(\theta)$ and $d(\pi-\theta)$, respectively. The sum and difference of their complex amplitudes, $M^{+}_{\bm{r}_{i}}\pm M^{-}_{\bm{r}_{i}}$, then form symmetric and antisymmetric combinations of these directional responses. For an ideal cardioid directivity, the symmetric (sum) component exhibits an omnidirectional response and thus provides an estimate of the local sound pressure. Conversely, the antisymmetric (difference) component approaches a figure-of-eight directivity aligned with $\bm{r}_{i}$, corresponding to the particle-velocity component along that axis.
Although practical cardioid microphones deviate from this ideal behavior, the pairwise sum and difference of oppositely oriented directional microphones still offer an effective operational approximation of the local pressure and velocity.
@@
$\widehat{p}_{\bm{r}_{i}}$ $=M^{+}_{\bm{r}_{i}}+M^{-}_{\bm{r}_{i}},$ (15)
$\widehat{u}_{\bm{r}_{i}}$ $=\frac{M^{+}_{\bm{r}_{i}}-M^{-}_{\bm{r}_{i}}}{Z_{0}},$ (16)
$\widehat{I}_{\bm{r}_{i}}$ $=\frac{|M^{+}_{\bm{r}_{i}}|^{2}-|M^{-}_{\bm{r}_{i}}|^{2}}{2Z_{0}},$ (17)
$\widehat{E}_{\bm{r}_{i}}$ $=\frac{|p_{\bm{r}_{i}}|^{2}}{4\rho_{0}c^{2}}+\frac{\rho_{0}}{4}|u_{\bm{r}_{i}}|^{2}.$ (18)
Figure 3: Schematic diagram of the signals used to construct the C-C method in the TF24 system. Twelve directional pairs are formed from 24 microphones.
Since the directional elements used in the array do not exhibit strictly cardioid directivity; their directivity varies with frequency. Therefore, the pseudo-pressure $\widehat{p}_{\bm{r}_{i}}$ obtained from their summation and the pseudo-velocity $\widehat{u}_{\bm{r}_{i}}$ inferred from their differential output cannot be regarded as the exact physical quantities. Instead, they represent operational approximations of the local sound pressure and particle velocity, derived under the assumption of complementary directivity pairs. The quantity corresponding to the intensity, $\widehat{I}_{\bm{r}_{i}}$, and the energy density, $\widehat{E}_{\bm{r}_{i}}$, are thus defined in this approximate sense and referred to here as pseudo-intensity and pseudo-energy density. Representative examples of these directivity patterns are presented in the next section.
Similarly, the directional dependent diffuseness index based on the I/E concept is defined as
$\Psi_{\bm{r}_{i}}=1-\frac{|\widehat{I}_{\bm{r}_{i}}|}{\>c\widehat{E}_{\bm{r}_{i}}\>},$ (19)
@@ III.1 Overview of simulation framework
The Afmt array was modeled after the AMBEO VR Mic, and the capsule directivities were approximated by a cardioid model consistent with DPA 2012. This microphone is also used for the TF24 array. The directivity patterns were measured in an anechoic chamber at $3^{\circ}$ resolution and fitted as a polynomial in $\cos\theta$ (power-series form) up to 8th order (as shown in Fig. 4); these fits were used in simulation.
Figure 4: Measured (solid lines) and modeled (broken lines) directivities of the DPA2012 cardioid microphone assumed in the Afmt and the TF24 array. To avoid visual clutter, the polar plots show the responses for 250 Hz to 1 kHz on the left half and for 2 kHz to 8 kHz on the right half. In the calculations presented in the text, however, a wider frequency range from 63 Hz to 16 kHz was analyzed in octave bands.
Figure 5: Representative example of modeled directional responses derived from the microphone directivity polynomial in Eq. (21) at 1 kHz. The results are shown (a) in linear magnitude and (b) in level (dB). The outer curve represents the pseudo-omnidirectional response obtained from the sum signal, whereas the inner curve corresponds to the pseudo-figure-of-eight response obtained from the difference signal. Both responses were derived from a pair of oppositely oriented microphones placed in the TF24 array with an inter-microphone spacing of 20 mm.
Array geometries were set as follows (see Fig. 2): the A-format capsules are located 6 mm from the origin [Fig. 2(a)]; the rigid-sphere array uses a radius of 42 mm [Fig. 2(b)]. The spacing between microphones in the TF24 array was set to 20 mm; that is, the distance from the origin in [Fig. 2(c)] is 10 mm. Although this dimension may seem unrealistic for practical microphones approximately 100 mm in length, it represents a practical configuration assuming sequential impulse-response measurements by mechanically rotating a single microphone.
@@
For the present simulations, the directivity pattern was modeled as a polynomial function of $\cos\theta$, where $\theta$ is the angle between $\hat{d}_{m}$ and $\hat{\nu}$:
@@
As an example, Fig. 5 illustrates the modeled directivities obtained from the sum and difference of opposite microphone pairs based on the polynomial model in Eq. (21) at 1 kHz. The sum response approximates a pseudo-omnidirectional pattern, while the difference response approximates a pseudo-figure-of-eight pattern.
For the Fibo64 rigid-sphere model with omnidirectional elements, $D_{m}=1$; directivity/baffle effects are captured by the radial filters $b_{n}(ka)$ in the Spherical Harmonic expansion. Thus, after computing $p_{m}(f)$ via (20), HOA coefficients are estimated from the SH model as described in Sec. II.
@@ III.2 Benchmark case definitions
A single plane wave arrives from a given direction; the theoretical diffuseness is $\Psi=0$. We sweep both elevation and azimuth over the sphere to test directional robustness and verify how accurately intensity-based direction-of-arrival can be recovered. For Case 1, 72 azimuthal directions (0 to 355) and 35 zenith angles (5 to 175), both at 5 intervals, were simulated, resulting in 2520 incidence directions in total. For each case, the mean diffuseness and the error in arrival-angle detection based on the pseudo-intensity were evaluated to assess the accuracy of sound-arrival direction estimation for each array.
@@
Case 1 and Case 2 are used not only as benchmark scenarios for diffuseness estimation but also as controlled tests to quantify the deviation between the proposed pseudo quantities and the ideal co-located pressure-velocity (C-C) reference.
@@
In this case, an interference field was simulated by superposing two plane waves: a primary wave incident from the zenith direction and a secondary wave arriving from variable zenith angles between 0° and 180°. For each configuration, 1,000 realizations were generated by randomly varying the phase (within $\pm 2\pi$) and amplitude (within $\pm 3$ dB) of the secondary wave. The pseudo-intensity, pseudo-energy density, and covariance matrix were computed for each realization, and the final values were obtained by averaging these quantities over all realizations to represent the mean behavior of the interference field.